Online Time Histories

Freyberger, Gilfoyle, Gyurjyan, and Wolin

Goal. Provide web-based tools to assess to the current CLAS data quality.

Status. The tools are working and in place. More input is needed from detector groups.

Time Histories Status

1. In place and working since January with online recsis and event monitor. Stadis added in April.

2. New features added and made more robust.

3. Demonstrated the capability of monitoring the current CLAS data quality.

4. Part of the e2 shift checklist.

To Be Done

1. Continue development.

(a) Add more time histories to online recsis, event monitor, and Stadis.

(b) Update online recsis to ET and improve speed.

(c) Add alarms!

2. GET MORE HELP/INPUT FROM THE DETECTOR GROUPS!

A Timeline History Primer

1. Generate your favorite monitoring variable in recsis.

2. Register the quantity in user init.F.

call timeline var reg(...)

3. Update the quantity in user evnt.F.

call timeline update db(...)

4. Monitor it through the experts time history web page. If you like it add it to the shift takers page